Debris Flow-2006/2007 Event Log

IOP-7

Lat: 34.66030 N            Lon: -118.76401 W          Alt: 814 m       Truck HD: 91 deg

Date/Time SR1 ready for operation: 19 March 2007 1710 UTC

 

Note taker: David Jorgensen (NSSL)

 

Time (UTC)

Event

19 March 1600

Arrived on site. Leveled truck. Booted computers. No problems

1615

Computer time off by ~5.5 hrs. Reset the time to match GPS handheld

1720

Radar up and running –no echoes anywhere – turning off transmitter – will take a scan every half hour or so to monitor the initiation of echoes

2055

Shutting down SR1 – will monitor the echoes from the hotel

20 March 1615

Arrived back on site. Updated forecast has echoes moving in this morning as front approached. Nothing missed last night.

1625

Leveled truck

1650

Radar up and running – convective looking echoes scattered about both north and south of burn area – max dbZ to about 45 or so

1712

Small echo over WHT

1750

Nav computer froze – rebooting

1753

Weak convective looking echoes over burn area max dbz to 40 or so

1754

Weak bright band on 90 deg scan – no drizzle at site yet

1757

A few raindrops now falling at site

1759

Convective echo near Buck Canyon USGS sites

1809

The few sprinkles of rain at site has subsided

1822

Another 45 dBZ echo over WTK

1918

Echoes around burn area decreasing in intensity

1930

Releveled truck – right rear leveler seems to sink a tad with time – backed up data to the laptop

1932

Only very small scattered echoes are left on the screen

1933

Going for lunch in Castaic

2058

Back from lunch – echoes about the same – some small weak storms over the burn area

2148

Some weak echoes over Buck and WTK areas

2212

Echoes growing again to the north of the burn area

2242

Rain falling at site again

2330

Going to dinner

0418

Back at SR1 – large echo over WTK max dbz on 8 deg tilt around 50 dbz

0440

Backup data to laptop

0450

Releveling truck. Lift front lifter down a few hundreds of a degree

0900

Data collection ended. Another meager rain event. WTK got only 0.14Ó from 2 PM – 10 PM PDT.